If the distance between two points is 55 mm on a 1:10 000 map, what would the distance in cm on a 1:50 000 map be?

1
Expert's answer
2021-08-20T01:27:02-0400

The distance will be smaller. Write a proportion:


"\frac{55}{x}=\frac{\frac{1}{10000}}{\frac{1}{50000}},\\\space\\
x=11\text{ mm}."